Happy Toes produces sports socks. The company has fixed expenses of $80,000 and variable expenses of $ 0.80 per package. Each package sells for $ 1.60

1.Provide the contribution margin formula. Compute the contribution margin per package and the contribution margin ratio. Enter to nearest cent.

2.Find the breakeven point in units and in dollars. Provide the formula and round to the nearest cent.

3.Find the number of packages Happy Toesneeds to sell to earn $25,000 operating income.
Answer #1

Happy Toes produces sports socks. The company has fixed expenses of $80,000 and variable expenses of $ 0.80 per package. Each package sells for $ 1.60

1.

Contribution margin per package = Selling price per package – Variable cost per package

= 1.60 - 0.80

= $0.80

Contribution margin ratio = Contribution margin per package/Selling price per package

= 0.80/1.60

= 50%

2.

Break even point (units) = Fixed cost/Contribution margin per package

= 80,000/0.80

= 100,000

Break even sales in dollars = Fixed cost/Contribution margin ratio

= 80,000/50%

= $160,000

3.

Packages to be sold to get a target profit = (Fixed cost + Target profit)/Contribution margin per package

= (80,000 + 25,000)/0.8

= 105,000/0.8

= 131,250